OpenStudy (anonymous):

find derivative of 4cos(2x)sin(y)=7

OpenStudy (anonymous):

I didn't get that result, but something similar.

OpenStudy (anonymous):

Let me know if you are still unclear.

OpenStudy (anonymous):

\[4\cos(2x)\sin(y)=7\\ \sin(y)= {7\over 4}\sec(2x) \\ \cos(y){dy\over dx}={7\over 4}\times2\times\tan(2x)\sec(2x)\\ {dy\over dx} ={7\over 2}\tan(2x)\sec(2x)\sec(y)\\ \]